OpenStudy (alfordabby):

can someone help me with my geometry homework? it's only three questions. I will fan & medal.

OpenStudy (anonymous):

Fire away and I will try to help! @alfordabby

OpenStudy (anonymous):

What is your question.

OpenStudy (alfordabby):

A shadow is cast from the building onto the ground. What is the height of the building if the shadow is 150 m?

OpenStudy (alfordabby):

@enderwither @tom_boy_rebel

OpenStudy (anonymous):

 a=b*tanA b=c*cosA a=c*sinA

OpenStudy (anonymous):

you are given A and a and you have to find b so since a=b*tanA, then b=a/tan A

OpenStudy (alfordabby):

this is what i have so far: xtan38 = 150/x xtan38/tan38 = 150/tan38 x=150/tan38 @ksanka

OpenStudy (alfordabby):

@Tom_Boy_Rebel

OpenStudy (alfordabby):

is that good so far? @ksanka